OpenStudy (tomektomeknyc):

sec (90 degrees - Arc Sin 1/2) = a) 1/2 b) 2 c)2 /√3 d)- 2

OpenStudy (anonymous):

\[\sec(\pi/2-\pi/6)=\sec(\pi/3)=1/\cos(\pi/3)=2\]

OpenStudy (tomektomeknyc):

Thank you